# Identification of Semiparametric Panel Multinomial

### v2

@@ A   Proof of Theorem 2
We first prove the convergence of $\hat{Q}_{j,t,s}\left(\beta\right)$ to $Q_{j,t,s}\left(\beta\right)$ for each $\left(j,t,s\right)$. For each generic deterministic function $\tilde{\gamma}_{j,t,s}$, define
$Q_{j,t,s}\left(\beta,\tilde{\gamma}\right)$ $:=\mathbb{E}\left[G\left(\tilde{\gamma}_{j,t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{j}\left({\bf X}_{i,ts};\beta\right)\right],$
$\hat{Q}_{j,t,s}\left(\beta,\tilde{\gamma}\right)$ $:=\frac{1}{n}\sum_{i=1}^{n}G\left(\tilde{\gamma}_{j,t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{j}\left({\bf X}_{i,ts};\beta\right).$
so that $\hat{Q}_{j,t,s}\left(\beta\right)=\hat{Q}_{j,t,s}\left(\beta,\tilde{\gamma}_{j,t,s}\right)$ and $Q_{j,t,s}\left(\beta\right)=Q_{j,t,s}\left(\beta,\gamma\right)$. For notational simplicity we suppress the subscript $\left(j,t,s\right)$ for the moment.
Defining ${\cal Q}:=\left\{G\left(\tilde{\gamma}\left(\overline{{\bf X}}\right)\right)\lambda\left({\bf X}_{i,ts};\beta\right):\tilde{\gamma}\in\Gamma,\beta\in\mathbb{\mathbb{S}}^{d-1}\right\}$, we first argue that ${\cal Q}$ is a $\mathbb{P}$-Donsker class based on Van Der Vaart and Wellner (1996). First, it is easy to show by Assumption 5 that $G\left(0\right)=0$, which together with the Lipschitz continuity of $G$, we have $\mathbb{E}\left[G^{2}\left(\tilde{\gamma}\left({\bf X}_{i}\right)\right)\right]\leq M\mathbb{E}\left[\tilde{\gamma}^{2}\left({\bf X}_{i}\right)\right]<\infty$ and $\mathbb{E}\left|G\left(\tilde{\gamma}\left({\bf X}_{i}\right)\right)\right|\leq\mathbb{E}\left|\tilde{\gamma}\left({\bf X}_{i}\right)\right|\leq\sup_{\tilde{\gamma}\in\Gamma}\mathbb{E}\left|\tilde{\gamma}\left({\bf X}_{i}\right)\right|<\infty$. Then, as $\Gamma$ is $\mathbb{P}$-Donsker, $G\circ\tilde{\gamma}$ must also be $\mathbb{P}$-Donsker. Second, recall that $\lambda\left({\bf X}_{i,ts};\beta\right)$ is the product of indicators of half planes, while the collection of $\mathbf{\mathbbm{1}}\left\{\left(\overline{X}_{k}-\underline{X}_{k}\right)^{{}^{\prime}}\beta\geq 0\right\}$ over $\beta\in\mathbb{\mathbb{S}}^{d-1}$ is a well-known VC Class of functions (sets) and is thus $\mathbb{P}$-Donsker. Finally, since the indicator function is uniformly bounded and $\sup_{\tilde{\gamma}\in\Gamma}\mathbb{E}\left|G\left(\tilde{\gamma}\left({\bf X}_{i}\right)\right)\right|<\infty$, we conclude that ${\cal Q}$ is also $\mathbb{P}$-Donsker:
@@
by Lipschitz continuity of $G$, $\left|\lambda_{j}\right|\leq 1$ and Cauchy-Schwarz. Combining (18) and (19), we have
$\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\left|\hat{Q}\left(\beta,\hat{\gamma}\right)-Q\left(\beta,\gamma\right)\right|\leq$ $\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\sup_{\tilde{\gamma}\in\Gamma}\left|\hat{Q}\left(\beta,\tilde{\gamma}\right)-Q\left(\beta,\tilde{\gamma}\right)\right|+\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\left|\hat{Q}\left(\beta,\hat{\gamma}\right)-Q\left(\beta,\hat{\gamma}\right)\right|$
@@
We verify Condition C.1 in Chernozhukov, Hong, and Tamer (2007, CHT thereafter) so as to apply their Theorem 3.1. Condition C.1(a) on the nonemptiness and compactness of parameter space is satisfied given Theorem 1. Condition C.1(b) on the continuity of the population criterion function $Q$ is proved by Lemma 3. Condition C.1(c) on measurability of the sample criterion function is satisfied by its construction. Condition C.1(d)(e) regarding the uniform convergence of $Q_{n}$ are satisfied by Lemma 4. Hence Theorem 3.1.(1) in CHT implies the Hausdorff consistency of $\hat{B}$. The consistency of the point estimator under the additional assumption of point identification (i.e., $B_{0}$ is a singleton) follows from Theorem 3.2 of CHT. ∎
Appendix B Pairwise Time Homogeneity of Errors
As mentioned in Section 2, Assumption 3 is stronger than necessary, and our identification strategy carries over under the weaker Assumption 3’, which requires that $\boldsymbol{\epsilon}_{it}\sim\left.\boldsymbol{\epsilon}_{is}\right|\left({\bf X}_{i,ts},{\bf A}_{i}\right).$ To see why Proposition 1 still holds, consider:
$\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right]$
$=$ $\int\mathbf{\mathbbm{1}}\left\{u\left(\overline{\delta}_{j},A_{ij},\epsilon_{ij{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}t}}\right)\geq\max_{k\neq j}u\left(\overline{\delta}_{k},\,A_{ik},\epsilon_{ik{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}t}}\right)\right\}\text{d}{\cal\mathbb{P}}\left(\left.\boldsymbol{\epsilon}_{i{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}t}}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right)$
$-\int\mathbf{\mathbbm{1}}\left\{u\left(\underline{\delta}_{j},A_{ij},\epsilon_{ij{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}}\right)\geq\max_{k\neq j}u\left(\underline{\delta}_{k},\,A_{ik},\epsilon_{ik{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}}\right)\right\}\text{d}{\cal\mathbb{P}}\left(\left.\boldsymbol{\epsilon}_{i{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right)$
$=$ $\int\mathbf{\mathbbm{1}}\left\{u\left(\overline{\delta}_{j},A_{ij},\tilde{\epsilon}_{ij}\right)\geq\max_{k\neq j}u\left(\overline{\delta}_{k},\,A_{ik},\tilde{\epsilon}_{ik}\right)\right\}\text{d}{\cal\mathbb{P}}\left(\left.\tilde{\boldsymbol{\epsilon}}_{i}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right)$
$-\int\mathbf{\mathbbm{1}}\left\{u\left(\underline{\delta}_{j},A_{ij},\tilde{\epsilon}_{ij}\right)\geq\max_{k\neq j}u\left(\underline{\delta}_{k},\,A_{ik},\tilde{\epsilon}_{ik}\right)\right\}\text{d}{\cal\mathbb{P}}\left(\left.\tilde{\boldsymbol{\epsilon}}_{i}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right)$
$=$ $\int\left[\begin{array}[]{c}\mathbf{\mathbbm{1}}\left\{u\left(\overline{\delta}_{j},A_{ij},\tilde{\epsilon}_{ij}\right)\geq\max_{k\neq j}u\left(\overline{\delta}_{k},\,A_{ik},\tilde{\epsilon}_{ik}\right)\right\}\\ -\mathbf{\mathbbm{1}}\left\{u\left(\underline{\delta}_{j},A_{ij},\tilde{\epsilon}_{ij}\right)\geq\max_{k\neq j}u\left(\underline{\delta}_{k},\,A_{ik},\tilde{\epsilon}_{ik}\right)\right\}\end{array}\right]\text{d}{\cal\mathbb{P}}\left(\left.\tilde{\boldsymbol{\epsilon}}_{i}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right)$
where $\overline{\boldsymbol{\delta}}=\overline{{\bf X}}\beta_{0}$, $\underline{\boldsymbol{\delta}}=\underline{{\bf X}}\beta_{0}$, and $\tilde{\boldsymbol{\epsilon}}_{i}$ denotes generic realizations of $\boldsymbol{\epsilon}_{it}$ and $\boldsymbol{\epsilon}_{is}$ conditional on ${\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ and ${\bf A}_{i}$. Notice that the second equality follows from the assumption that $\boldsymbol{\epsilon}_{it}\sim\left.\boldsymbol{\epsilon}_{is}\right|\left({\bf X}_{i,ts},{\bf A}_{i}\right)$.
Again, if $\overline{\delta}_{j}\leq\underline{\delta}_{j}\text{ and }\overline{\delta}_{k}\geq\underline{\delta}_{k}\text{ for all }k\neq j,$ the bracketed term in the last line of the displayed equation above must be nonpositive for all realizations of ${\bf A}_{i}$ and $\tilde{\boldsymbol{\epsilon}}_{i}$, so that $\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{it}=\overline{{\bf X}},{\bf X}_{is}=\underline{{\bf X}},{\bf A}_{i}\right]\leq 0$ for all realizations of ${\bf A}_{i}$, which further implies that
$\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)=\mathbb{E}\left[\left.\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right]\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right]\leq 0.$
Taking the logical contraposition again gives Proposition 1.
@@ Appendix C Sufficient Conditions for Point Identification
In this section, we prove sufficient conditions for the point identification of $\beta_{0}$. For simplicity of notation, we fix $T=2$. We first need to impose an assumption of strict multivariate monotonicity on the function $\psi_{j}$ defined in (4).
@@
For any realized ${\bf A}_{i}$, the function $\psi_{j}\left(\,\boldsymbol{\cdot}\,,{\bf A}_{i}\right):\mathbb{R}^{J}\to\mathbb{R}$ is strictly increasing, i.e., if $\overline{\delta}_{j}>\underline{\delta}_{j}$ for all $j$, then $\psi\left(\overline{\boldsymbol{\delta}},{\bf A}_{i}\right)>\psi\left(\underline{\boldsymbol{\delta}},{\bf A}_{i}\right).$
@@
Assumption 1’ (Strict Monotonicity of $u$). $u\left(\delta_{ijt},A_{ij},\epsilon_{ijt}\right)$ is strictly increasing in the index $\delta_{ijt}$, for every realization of $\left(A_{ij},\epsilon_{ijt}\right)$.
Assumption 1” (Overlapping Supports). Conditional on any realization of ${\bf X}_{i}$ and ${\bf A}_{i}$, we have $\bigcap_{j=1}^{J}int\left(Supp\left(u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)\right)\right)\neq\emptyset$.
In particular, Assumption LABEL:as1pp is directly implied by the assumption of $Supp\left(u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)\right)=\mathbb{R}$ conditional on any realization of ${\bf X}_{i}$ and ${\bf A}_{i}$, which is again satisfied in additive panel multinomial choice models with scalar fixed effects a la $u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)=X_{ijt}^{{}^{\prime}}\beta_{0}+A_{ij}+\epsilon_{ijt}$ under the assumption of $Supp\left(\left.\epsilon_{ijt}\right|{\bf X}_{i},{\bf A}_{i}\right)=\mathbb{R}$ as commonly imposed in the literature.
@@
Finally, we impose the following assumption on $\Delta{\bf X}_{i}$, with $\Delta X_{ij}:=X_{ij1}-X_{ij2}$ for all individual $i$ and product $j$ across period 1 and period 2.
@@
    There exists some $k\in\left\{1,...,d_{x}\right\}$ such that $\beta_{0}^{k}\neq 0$ and $Supp\left(\left.\Delta X_{ij}^{k}\right|\Delta X_{il},\,l\neq j\right)=\mathbb{R}$ for all $j\in\left\{1,...,J\right\}$. Furthermore, for all $j\in\left\{1,...,J\right\}$, $Supp\left(\left.\Delta X_{ij}\right|\Delta X_{il},\,l\neq j\right)$ is not contained in a proper linear subspace of $\mathbb{R}^{d_{x}}$.
Assumption 8(a) is satisfied when $\left(X_{ij}\right)$ is continuous random vector. On the other hand, Assumption 8(b) can accommodate discrete regressors generally, but requires one continuous covariate with large support. Assumption 8 ensures that $\Delta X_{ij}^{{}^{\prime}}\beta_{0}>0$ and $\Delta X_{ik}^{{}^{\prime}}\beta_{0}<0$ for all $k\neq j$ hold simultaneously with strictly positive probability.
@@
Under Assumptions 1, 3, 7 and 8, $\beta_{0}$ is point identified on $\mathbb{\mathbb{S}}^{D-1}$.
@@
Recall first that
$\gamma_{j}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ $=\int\left[\psi_{j}\left(\overline{\delta}_{j},\left(-\overline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)-\psi_{j}\left(\underline{\delta}_{j},\left(-\underline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)\right]\text{d}\mathbb{P}\left(\left.{\bf A}_{i}\right|{\bf X}_{i}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right).$
@@
$\overline{\delta}_{j}<\underline{\delta}_{j}\text{ and }\overline{\delta}_{k}>\underline{\delta}_{k}\text{ for all }k\neq j\quad\Rightarrow\quad\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0,$ (20)
since $\psi_{j}\left(\overline{\delta}_{j},\left(-\overline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)<\psi_{j}\left(\underline{\delta}_{j},\left(-\underline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)$ for every realization of ${\bf A}_{i}$. Together with Assumption 8, we deduce that
@@
Now for any $\beta\in\mathbb{\mathbb{S}}^{D-1}\backslash\left\{\beta_{0}\right\}$, define for any product $j$,
@@
As $\beta\neq\beta_{0}$, by Assumption 8 we know that
@@
Moreover, for any realization of ${\bf X}_{i}$ s.t. $\Delta{\bf X}_{i}\in H_{j}\left(\beta\right)$, we must have: (i) $\gamma_{j,t,s}\left({\bf X}_{i}\right)>0$ by (20), and (ii):
@@
so that $G\left(\gamma_{j}\left({\bf X}_{i}\right)\right)\lambda_{j}\left(\Delta{\bf X}_{i},\beta\right)=G\left(\gamma_{j}\left({\bf X}_{i}\right)\right)>0$ for all such ${\bf X}_{i}$. Hence,

### v3

@@ Appendix B Proof of Theorem 3
We first prove the convergence of $\hat{Q}_{\tilde{{\cal J}},t,s}\left(\beta\right)$ to $Q_{\tilde{{\cal J}},t,s}\left(\beta\right)$ for each $\left(\tilde{{\cal J}},t,s\right)$. For each generic deterministic function $\tilde{\gamma}_{\tilde{{\cal J}},t,s}$, define
[…]
so that $\hat{Q}_{\tilde{{\cal J}},t,s}\left(\beta\right)=\hat{Q}_{\tilde{{\cal J}},t,s}\left(\beta,\tilde{\gamma}_{\tilde{{\cal J}},t,s}\right)$ and $Q_{\tilde{{\cal J}},t,s}\left(\beta\right)=Q_{\tilde{{\cal J}},t,s}\left(\beta,\gamma\right)$. For notational simplicity, we suppress the subscript $\left(\tilde{{\cal J}},t,s\right)$ for the moment.
Defining ${\cal Q}:=\left\{G\left(\tilde{\gamma}\left({\bf x}_{ts}\right)\right)\lambda\left({\bf x}_{ts};\beta\right):\tilde{\gamma}\in\Gamma,\beta\in\mathbb{\mathbb{S}}^{d-1}\right\}$, we first argue that $\cal Q$} is a $\mathbb{P}$-Donsker class based on Van Der Vaart and Wellner (1996). First, it is easy to show by Assumption 5 that $G\left(0\right)=0$, which together with the Lipschitz continuity of $G$, implies that $\mathbb{E}\left[G^{2}\left(\tilde{\gamma}\left({\bf X}_{i}\right)\right)\right]\leq M\mathbb{E}\left[\tilde{\gamma}^{2}\left({\bf X}_{i}\right)\right]<\infty$ and $\mathbb{E}\left|G\left(\tilde{\gamma}\left({\bf X}_{i}\right)\right)\right|\leq\mathbb{E}\left|\tilde{\gamma}\left({\bf X}_{i}\right)\right|\leq\sup_{\tilde{\gamma}\in\Gamma}\mathbb{E}\left|\tilde{\gamma}\left({\bf X}_{i}\right)\right|<\infty$. Then, as $\Gamma$ is $\mathbb{P}$-Donsker, $G\circ\tilde{\gamma}$ must also be $\mathbb{P}$-Donsker. Second, recall that $\lambda\left({\bf X}_{i,ts};\beta\right)$ is the product of indicators of half planes, while the collection of $\mathbf{\mathbbm{1}}\left\{\left({\bf x}_{kt}-{\bf x}_{ks}\right)^{{}^{\prime}}\beta\geq 0\right\}$ over $\beta\in\mathbb{\mathbb{S}}^{d-1}$ is a well-known VC-class of functions (sets) and is thus $\mathbb{P}$-Donsker. Finally, since the indicator function is uniformly bounded and $\sup_{\tilde{\gamma}\in\Gamma}\mathbb{E}\left|G\left(\tilde{\gamma}\left({\bf X}_{i}\right)\right)\right|<\infty$, we conclude that $\cal Q$} is also $\mathbb{P}$-Donsker:
@@
$\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\left|Q\left(\beta,\hat{\gamma}\right)-Q\left(\beta,\gamma\right)\right|\leq$ $\sup_{\beta\in\mathbb{\mathbb{S}}^{d-1}}\int\left|G\left(\hat{\gamma}\left({\bf x}_{ts}\right)\right)-G\left(\gamma\left({\bf x}_{ts}\right)\right)\right|\lambda_{j}\left({\bf x}_{ts};\beta\right)d\mathbb{P}\left({\bf x}_{ts}\right)$
$\leq$ $\ M\sqrt{\int\left(\hat{\gamma}\left({\bf x}_{ts}\right)-\gamma\left({\bf x}_{ts}\right)\right)^{2}d\mathbb{P}\left({\bf x}_{ts}\right)}=O_{p}\left(c_{N}\right)$ (39)
by Lipschitz continuity of $G$, $\left|\lambda_{j}\right|\leq 1$, and the Cauchy-Schwarz inequality. Combining (38) and (39), we have
@@
We verify Condition C.1 in Chernozhukov, Hong, and Tamer (2007, CHT thereafter) so as to apply their Theorem 3.1. Condition C.1(a) on the non-emptiness and compactness of parameter space is satisfied given Theorem 1. Condition C.1(b) on the continuity of the population criterion function $Q$ is proved by Lemma 2. Condition C.1(c) on measurability of the sample criterion function is satisfied by construction. Conditions C.1(d)(e) regarding the uniform convergence of $Q_{n}$ are satisfied by Lemma 3. Hence, Theorem 3.1.(1) in CHT implies the Hausdorff consistency of $\hat{B}$. The consistency of the point estimator under the additional assumption of point identification (i.e., $B_{0}$ is a singleton) follows from Theorem 3.2 of CHT. ∎
@@ Appendix C Sufficient Conditions for Point Identification
In this section, we prove sufficient conditions for the point identification of $\beta_{0}$. For simplicity of notation, we fix $T=2$, and focus on the conditions that would establish point identification using identifying restrictions from singleton products $j=1,\ldots,J$. Consequently, these sufficient conditions should be regarded as being strictly more than necessary.
Define $\boldsymbol{\delta}_{t}:=\left(\delta_{1t},\ldots,\delta_{Jt}\right)^{{}^{\prime}}$, where $\delta_{jt}:=x_{jt}^{{}^{\prime}}\beta_{0}$, and recall that $\delta_{ijt}=X_{ijt}^{{}^{\prime}}\beta_{0}$. Denote
[…]
We first impose a strict multivariate monotonicity condition on $\psi_{j}\left(\cdot\right)$.
@@
For any realized ${\bf A}_{i}$ and ${\bf x}_{ts}$, the function $\psi_{j}\left(\boldsymbol{\delta}_{t};{\bf x}_{ts},{\bf A}_{i}\right):\mathbb{R}^{J}\to\mathbb{R}$ is strictly increasing in $\delta_{jt}$ and decreasing in $\delta_{kt}$ for $k\neq j$, i.e., if for any two periods $t$ and $s$ it is true that $\delta_{jt}>\delta_{js}$ and $\delta_{kt}<\delta_{ks}$ for all $k\neq j$, then $\psi_{j}\left(\boldsymbol{\delta}_{t};{\bf x}_{ts},{\bf A}_{i}\right)>\psi_{j}\left(\boldsymbol{\delta}_{s};{\bf x}_{ts},{\bf A}_{i}\right).$
@@
Assumption 2 (Strict Monotonicity of $u$). $u\left(\delta_{ijt},A_{ij},\epsilon_{ijt}\right)$ is strictly increasing in the index $\delta_{ijt}$, for every realization of $\left(A_{ij},\epsilon_{ijt}\right)$.
Assumption 2′′ (Overlapping Supports). Conditional on any realization of ${\bf X}_{i}$ and ${\bf A}_{i}$, we have $\bigcap_{j=1}^{J}int\left(Supp\left(u\left(\delta_{ijt},A_{ij},\epsilon_{ijt}\right)\right)\right)\neq\emptyset$.
In particular, Assumption LABEL:as1pp is directly implied by the assumption of $\text{Supp}\left(u\left(\delta_{ijt},A_{ij},\epsilon_{ijt}\right)\right)$ being equal to the real line conditional on any realization of ${\bf X}_{i}$ and ${\bf A}_{i}$, which is again satisfied in additive panel multinomial choice models with scalar fixed effects a la
$u\left(\delta_{ijt},A_{ij},\epsilon_{ijt}\right)=\delta_{ijt}+A_{ij}+\epsilon_{ijt}$
under the assumption of $\text{Supp}\left(\left.\epsilon_{ijt}\right|{\bf X}_{i},{\bf A}_{i}\right)=\mathbb{R}$ as commonly imposed in the literature.
@@
Finally, we impose the following assumption on $\Delta{\bf X}_{i}$, with $\Delta X_{ij}:=X_{ij1}-X_{ij2}$ for all individual $i$ and product $j$ between period 1 and period 2.
@@
    There exists some $k\in\left\{1,\ldots,D\right\}$ such that $\beta_{0}^{k}\neq 0$ and $Supp\left(\left.\Delta X_{ij}^{k}\right|\Delta X_{il},\,l\neq j\right)=\mathbb{R}$ for all $j\in\left\{1,\ldots,J\right\}$. Furthermore, for all $j\in\left\{1,\ldots,J\right\}$, $Supp\left(\left.\Delta X_{ij}\right|\Delta X_{il},\,l\neq j\right)$ is not contained in a proper linear subspace of $\mathbb{R}^{D}$.
Assumption 8(a) is satisfied when $X_{ij}$ is a continuous random vector. On the other hand, Assumption 8(b) can accommodate discrete regressors generally, but requires one continuous covariate with a large support. Assumption 8 ensures that $\Delta X_{ij}^{{}^{\prime}}\beta_{0}>0$ and $\Delta X_{ik}^{{}^{\prime}}\beta_{0}<0$ for all $k\neq j$ hold simultaneously with strictly positive probability.
@@
Under Assumptions 1, 3, 7, and 8, $\beta_{0}$ is point identified on $\mathbb{\mathbb{S}}^{D-1}$.
@@
Fix any ${\bf x}_{ts}$ in the support of ${\bf X}_{i}$. By definition of $\gamma_{j,t,s}$, we have
$\gamma_{j,t,s}\left({\bf x}_{ts}\right)$ $=\int\left[\psi_{j}\left(\boldsymbol{\delta}_{t};{\bf x}_{ts},{\bf A}_{i}\right)-\psi_{j}\left(\boldsymbol{\delta}_{s};{\bf x}_{ts},{\bf A}_{i}\right)\right]\text{d}\mathbb{P}\left(\left.{\bf A}_{i}\right|{\bf X}_{i,ts}={\bf x}_{ts}\right).$
@@
$\delta_{jt}>\delta_{js}\text{ and }\delta_{kt}<\delta_{ks}\text{ for all }k\neq j\quad\Rightarrow\quad\gamma_{j,t,s}\left({\bf x}_{ts}\right)>0,$ (40)
since $\psi_{j}\left(\boldsymbol{\delta}_{t};{\bf x}_{ts},{\bf A}_{i}\right)>\psi_{j}\left(\boldsymbol{\delta}_{s};{\bf x}_{ts},{\bf A}_{i}\right)$ for every realization of ${\bf A}_{i}$ given ${\bf X}_{i,ts}={\bf x}_{ts}$. Together with Assumption 8, we deduce that
@@
[…]
@@
As $\beta\neq\beta_{0}$, by Assumption 8 we have
@@
Moreover, for any realization of ${\bf X}_{i}$ such that $\Delta{\bf X}_{i}\in H_{j}\left(\beta\right)$, we must have: (i) $\gamma_{j,t,s}\left({\bf X}_{i}\right)>0$ by (40), and (ii)
@@
so that
$G\left(\gamma_{j}\left({\bf X}_{i}\right)\right)\lambda_{j}\left(\Delta{\bf X}_{i},\beta\right)=G\left(\gamma_{j}\left({\bf X}_{i}\right)\right)>0$
for all such ${\bf X}_{i}$. Hence, we have